OpenStudy (anonymous):

How do I solve a problem with 3 radicals.

OpenStudy (anonymous):

\[\sqrt{x+\sqrt{2x}} = \sqrt{2x}\]

ganeshie8 (ganeshie8):

start by squaring both sides

OpenStudy (anonymous):

\[(\sqrt{x+\sqrt{2x}})^{2} = (\sqrt{2x})^{2}\] like this?

ganeshie8 (ganeshie8):

yes cancel the radical

OpenStudy (anonymous):

\[x+\sqrt{2x} = 2x\] Now what do I do?

ganeshie8 (ganeshie8):

subtract x both sides

OpenStudy (anonymous):

Would it be\[\sqrt{2x} = 2x-x\]

OpenStudy (anonymous):

To give me \[\sqrt{2x} = x\]

ganeshie8 (ganeshie8):

yes ! square it again

OpenStudy (anonymous):

Then I would have,\[2x = x^{2}\] is this true?

ganeshie8 (ganeshie8):

Correct !

ganeshie8 (ganeshie8):

its a quadratic, u can factor it and solve x